Last month a man paid $65 for 5 dvds and 2 cd's. This month he paid $81 for 3 dvd's and 4 cd's. How much were the DVD's and CD's?
Answer by mananth(16949)

let the dvd cost x
let the cd cost y
last month
5x+2y=65
this month
3x+4y =81
3(5x+2y)= 3*65
5(3x+4y)=5*81
18x+6y = 195
18x+20y=405
subtract the equations
-14y=- 210
y=-210/-14
y=15
By substituting the value of y in any equation we get x=7


x=7